[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Dec 15, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Dec 15 15:42:58 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
No new issues.
HSU:
No new issues.
Other items:
1. This morning's 100G upgrade of the network at the WCOSS facility in  
Orlando was "more impactive than anticipated." All development queues have  
been closed (for approximately 30 minutes) and the shared filesystem (/sss)  
has been unmounted.

2. At 0359Z the /prod00/wave/multi_1/jwave_multi_1_forecast (global  
multi-grid WaveWatch III model) job failed with a severity level 1 error.  
After the winds were updated, the job was rerun to completion at 0517Z.  
Operational Impact: 00Z Wave multi_1 forecast up to 43 minutes late.

3. Slow GFS dissemination via NCF: Beginning at 1623Z yesterday, many  
offices reported missing GFS model forecast fields and slow  
dissemination/loading into AWIPS. This was due to an NCF socket test. At  
1705Z, the queue of delayed products began bleeding down.

4. GOES-EAST satellite remap server: At 1821Z yesterday, the geodist2  
(GOES-13/E remapper) was reset, the GOES-E ingester was unable to  
communicate with it, and imagery was not updating. At 2115Z, the problem  
was resolved and images were ingesting properly. No impact within NCEP...  
all stitched GOES-E/W products looked okay in the SDM's AWIPS and NAWIPS.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161215/1514d376/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list